State the range of function f(x)=(4+x3)2 with the given domain -1<x≤1 where x∈ℝ.

1
Expert's answer
2021-10-01T04:24:32-0400

Range is the value of f(x) for the given set of input which is the value of x. When x=-1, f(x)=(4+x3)2

f(-1)=(4-1)2=9

f(1)=(4+1)2=25

Thus range is from 9 to 25